OpenStudy (anonymous):

Solve Each Equation by first clearing the fractions or decimals. Show all work 1/3-1/2m=1/6m+7/9

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }-\frac{ 1 }{ 2 }m=\frac{ 1 }{ 6 }m +\frac{ 7 }{ 9 }\]

OpenStudy (wikiemol):

add (1/2)m to both sides 1/3 = 1/6m + 1/2m + 7/9 subtract 7/9 from both sides 1/3 - 7/9 = 1/6m + 1/2m make the denominators the same 3/9 - 7/9 = 1/6m + 3/6m add and subtract -4/9 = 4/6m divide 4/6 from both sides (the same as multiplying each side by the reciprocal) (-4/9)*(6/4) = m -24/36 = m simplify -2/3 = m